# XAI Evaluation Framework for Semantic Segmentation

### v1

@@ Abstract
Ensuring transparency and trust in artificial intelligence (AI) models is essential, particularly as they are increasingly applied in safety-critical and high-stakes domains. Explainable AI (XAI) has emerged as a promising approach to address this challenge, yet the rigorous evaluation of XAI methods remains crucial for optimizing the trade-offs between model complexity, predictive performance, and interpretability. While extensive progress has been achieved in evaluating XAI techniques for classification tasks, evaluation strategies tailored to semantic segmentation remain relatively underexplored. This work introduces a comprehensive and systematic evaluation framework specifically designed for assessing XAI in semantic segmentation, explicitly accounting for both spatial and contextual task complexities. The framework employs pixel-level evaluation strategies and carefully designed metrics to provide fine-grained interpretability insights. Simulation results using recently adapted class activation mapping (CAM)-based XAI schemes demonstrate the efficiency, robustness, and reliability of the proposed methodology. These findings contribute to advancing transparent, trustworthy, and accountable semantic segmentation models.
@@ 3 Proposed XAI Evaluation Framework
XAI U PM: where we combine the XAI heatmap with the predicted target areas of the model to see which XAI method highlights the most relevant pixels for model prediction. This allows us to visualize which XAI method shows the most important pixels that help the model make its predictions, even if those predictions are not perfect.
XAI U GT: comparing XAI-highlighted pixels with actual GT data to assess which method accurately identifies pixels that truly contribute to optimizing the model performance. This helps to know which XAI method highlights the pixels that actually help to improve the prediction of the model. The XAI heatmap is wrapped over the ground truth (GT) target mask. The edited image displays both the real target class pixels from the GT mask and the ones that have been highlighted. In this manner, we are able to measure how selective each XAI method is in highlighting those specific pixels that actually would contribute to optimizing the model’s performance.
@@ 4.2 S2: highlighted only
4.3 S3: XAI U GT
In the context of XAI methods for segmentation tasks, highlighting inside and outside the target class refers to how well these methods identify not only the pixels that belong to the target class but also those that might not directly belong to the target class but still play a significant role in the model’s decision-making process. To assess the sensitivity of XAI methods on segmenting all the target pixels, we compare the union of the highlighted pixels generated by the XAI and the ground truth segmentation mask (XAI U GT). This measure identifies to what degree the explanation maps include relevant target class pixels and potentially explanatory neighboring pixels. It provides a baseline for understanding whether the XAI methodology picks up all semantically important regions as defined by the ground truth to set the scene for comparison to the model’s own predicted outcome in the subsequent section.
@@ 4.3 S3: XAI U GT
4.4 S3: XAI U PM
To further validate the insights derived from the XAI U GT evaluation, we also performed an analysis using the predicted segmentation masks (XAI U PM). This comparison assesses how well the explanation maps align with the model’s own output, offering a practical perspective on the interpretability of the segmentation predictions. Interestingly, the results for XAI U PM closely mirror those obtained from XAI U GT in terms of TP, FP, precision, recall, and IoU. This consistency reinforces the reliability of the XAI methods and confirms that the highlighted pixels, whether aligned with the ground truth or the model’s prediction, capture relevant information that influences the model’s decision-making process. It also emphasizes that surrounding contextual pixels remain essential, regardless of the reference mask used, further validating the conclusions drawn in the previous section.
As shown in Tables 7 and 8, the results of XAI U PM align with the findings of XAI U GT, supporting the consistency and reliability of XAI methods. Score-CAM continues to over perform, achieving the highest TP pixel count (47,391), precision (0.47), recall (0.89), and IoU (0.45), indicating its ability to accurately highlight relevant regions that align closely with the model’s output. Eigen-CAM also shows balanced performance with high TP counts, a moderate FP rate, and an IoU of 0.41, confirming its ability to capture both central and contextual pixels effectively. Grad-CAM++ maintains its improved focus compared to Grad-CAM, with better precision (0.35) and recall (0.88), while Grad-CAM and Ablation-CAM still maintain higher FP rates and lower IoU scores, suggesting overgeneralized highlighting. XGrad-CAM still reflects moderate performance. Overall, Score-CAM and Eigen-CAM remain the most interpretable and reliable methods in terms of alignment with the model’s predictions, while others demonstrate trade-offs between recall and precision. These findings confirm that the highlighted regions, whether they intersect with ground truth or the prediction of the model, carry meaningful information and that context beyond the pixels of the target class significantly affects the model’s decision process.

### v2

# A Quantitative Evaluation Framework for Explainable AI in Semantic Segmentation

@@ Abstract
Ensuring transparency and trust in artificial intelligence (AI) models is essential as they are increasingly deployed in safety-critical and high-stakes domains. Explainable AI (XAI) has emerged as a promising approach to address this challenge; however, the rigorous evaluation of XAI methods remains vital for balancing the trade-offs between model complexity, predictive performance, and interpretability. While substantial progress has been made in evaluating XAI for classification tasks, strategies tailored to semantic segmentation remain limited. Moreover, objectively assessing XAI approaches is difficult, since qualitative visual explanations provide only preliminary insights. Such qualitative methods are inherently subjective and cannot ensure the accuracy or stability of explanations. To address these limitations, this work introduces a comprehensive quantitative evaluation framework for assessing XAI in semantic segmentation, accounting for both spatial and contextual task complexities. The framework systematically integrates pixel-level evaluation strategies with carefully designed metrics to yield fine-grained interpretability insights. Simulation results using recently adapted class activation mapping (CAM)-based XAI schemes demonstrate the efficiency, robustness, and reliability of the proposed methodology. These findings advance the development of transparent, trustworthy, and accountable semantic segmentation models.
@@ 3 Proposed XAI Evaluation Framework
XAI-PM: where we combine the XAI heatmap with the predicted target areas of the model to see which XAI method highlights the most relevant pixels for model prediction. This allows us to visualize which XAI method shows the most important pixels that help the model make its predictions, even if those predictions are not perfect.
XAI-GT: comparing XAI-highlighted pixels with actual GT data to assess which method accurately identifies pixels that truly contribute to optimizing the model performance. This helps to know which XAI method highlights the pixels that actually help to improve the prediction of the model. The XAI heatmap is wrapped over the ground truth (GT) target mask. The edited image displays both the real target class pixels from the GT mask and the ones that have been highlighted. In this manner, we are able to measure how selective each XAI method is in highlighting those specific pixels that actually would contribute to optimizing the model’s performance.
@@ 4.2 S2: highlighted only
4.3 S3: XAI-GT
In the context of XAI methods for segmentation tasks, highlighting inside and outside the target class refers to how well these methods identify not only the pixels that belong to the target class but also those that might not directly belong to the target class but still play a significant role in the model’s decision-making process. To assess the sensitivity of XAI methods on segmenting all the target pixels, we compare the union of the highlighted pixels generated by the XAI and the ground truth segmentation mask (XAI-GT). This measure identifies to what degree the explanation maps include relevant target class pixels and potentially explanatory neighboring pixels. It provides a baseline for understanding whether the XAI methodology picks up all semantically important regions as defined by the ground truth to set the scene for comparison to the model’s own predicted outcome in the subsequent section.
@@ 4.3 S3: XAI-GT
4.4 S3: XAI-PM
To further validate the insights derived from the XAI-GT evaluation, we also performed an analysis using the predicted segmentation masks (XAI-PM). This comparison assesses how well the explanation maps align with the model’s own output, offering a practical perspective on the interpretability of the segmentation predictions. Interestingly, the results for XAI-PM closely mirror those obtained from XAI-GT in terms of TP, FP, precision, recall, and IoU. This consistency reinforces the reliability of the XAI methods and confirms that the highlighted pixels, whether aligned with the ground truth or the model’s prediction, capture relevant information that influences the model’s decision-making process. It also emphasizes that surrounding contextual pixels remain essential, regardless of the reference mask used, further validating the conclusions drawn in the previous section.
As shown in Tables 7 and 8, the results of XAI-PM align with the findings of XAI-GT, supporting the consistency and reliability of XAI methods. Score-CAM continues to over perform, achieving the highest TP pixel count (47,391), precision (0.47), recall (0.89), and IoU (0.45), indicating its ability to accurately highlight relevant regions that align closely with the model’s output. Eigen-CAM also shows balanced performance with high TP counts, a moderate FP rate, and an IoU of 0.41, confirming its ability to capture both central and contextual pixels effectively. Grad-CAM++ maintains its improved focus compared to Grad-CAM, with better precision (0.35) and recall (0.88), while Grad-CAM and Ablation-CAM still maintain higher FP rates and lower IoU scores, suggesting overgeneralized highlighting. XGrad-CAM still reflects moderate performance. Overall, Score-CAM and Eigen-CAM remain the most interpretable and reliable methods in terms of alignment with the model’s predictions, while others demonstrate trade-offs between recall and precision. These findings confirm that the highlighted regions, whether they intersect with ground truth or the prediction of the model, carry meaningful information and that context beyond the pixels of the target class significantly affects the model’s decision process.
Finally, we would like to mention that the proposed Quantitative XAI Evaluation Framework is not limited to remote sensing applications. It is built on a generic methodology that addresses: (i) Fidelity, i.e., does the explanation truly reflect the model’s decision-making strategy?, and (ii) robustness, i.e., are the explanations stable and reliable? Hence, the proposed XAI evaluation framework could be easily applied to a wide range of semantic segmentation applications. For example, in AI-based medical imaging [Muhammad and Bendechache, 2024], it provides the objective metrics to ensure that the related XAI methods are clinically sound. In short, our work advances reliable XAI for any critical application requiring trustworthy AI-based semantic segmentation.